# Testing Monotonicity in a Finite Population

### v1

@@ 5. Bayesian updating
    Some Bayesians update: there exists a prior $\pi$ with $\pi(\Theta_{0})\in(0,1)$ such that $\pi(\Theta_{0}\mid Y)\neq\pi(\Theta_{0})$ with positive $\pi$-probability.
@@
    There exist (nontrivial) Bayesian priors over $\theta$ that never update about the probability that monotonicity holds: for any $c\in(0,1)$, there exists a prior distribution $\pi$ over $\theta$ such that $\pi(\theta\in\Theta_{0}\mid Y)=\pi(\theta\in\Theta_{0})=c$, $\pi$-almost surely.
@@ 6. Conclusion
We study what a completely randomized experiment can reveal about finite-population monotonicity. We show that from the design-based perspective, the type shares $\theta$ in the finite population are in fact identified. However, the extent to which we can feasibly learn about violations of monotonicity is severely limited: frequentist tests generically have poor power, and some Bayesians never update about whether the null is true. Thus, formal identification translates to little, if any, practical learning about monotonicity. These results highlight that conclusions about identification may differ depending on whether one adopts a sampling-based versus design-based perspective, and that studying the properties of frequentist tests and Bayesian updating may provide a more realistic assessment of the extent to which learning is possible in design-based settings. An interesting avenue for future research is to explore whether similar issues arise in other design-based causal inference problems.
@@ Appendix A Proofs
For the first claim, it suffices to construct a mapping from $f_{\theta}$ to $\theta$. Fix $q=n_{1}/n$ as the treatment probability. Observer that, by linearity of expectation,
@@
Second, label underlying units $i=1,\ldots,n$ and collect always takers in $N_{a}\subset[n]$, compliers in $N_{c}\subset[n]$, and defiers in $N_{d}\subset[n]$. Let $N_{1}\subset[n]$ collect the (random) set of treated individuals. Observe that
$E_{\theta}[Y_{T}Y_{U}]$ $=\sum_{i\neq j,i\in N_{a}\cup N_{c},j\in N_{a}\cup N_{d}}P(\text{$i\in N_{1}$, $j\not\in N_{1}$})$ (Linearity of expectations)
$=|\{(i,j)\,:\,i\neq j,i\in N_{a}\cup N_{c},j\in N_{a}\cup N_{d}\}|\cdot P(\text{$1\in N_{1}$, $2\not\in N_{1}$})$ (Symmetry)
$=\left[{(n_{at}+n_{c})(n_{at}+n_{d})-n_{a}}\right]P(\text{$1\in N_{1}$, $2\not\in N_{1}$}).$
Thus,
$(n_{at}+n_{c})(n_{at}+n_{d})-n_{at}=\frac{E_{\theta}[Y_{T}Y_{U}]}{P(\text{$1\in N_{1}$, $2\not\in N_{1}$})}$
is a function of $f_{\theta}$. Letting $(\mu_{T},\mu_{Y},\mu_{TU}):=(E_{\theta}[Y_{T}],E_{\theta}[Y_{U}],E_{\theta}[Y_{T}Y_{U}])$, we see that
$\frac{\mu_{T}}{q}=n_{at}+n_{c},\hskip 28.45274pt\frac{\mu_{U}}{1-q}=n_{at}+n_{d},\hskip 28.45274pt\mu_{TU}=\frac{\mu_{T}}{q}\,\frac{\mu_{U}}{1-q}-n_{at}.$
@@
$n_{at}=\mu_{TU}-\frac{\mu_{T}}{q}\,\frac{\mu_{U}}{1-q}\hskip 28.45274ptn_{c}=\frac{\mu_{T}}{q}-n_{at},\hskip 28.45274ptn_{d}=\frac{\mu_{U}}{1-q}-n_{at}.$
Since $(\mu_{T},\mu_{Y},\mu_{TU})$ is a function of $f_{\theta}$, we have therefore shown that there is an inverse mapping from $f_{\theta}$ to $\theta$, and hence the mapping $\theta\mapsto f_{\theta}$ must be one-to-one, as needed.
@@
Lemma A.1.
Suppose $n_{1}=\frac{n}{2}=n_{0}$. Then there exist priors $\pi_{A}$ and $\pi_{B}$ on $\theta$ such that $\pi_{A}(\theta\in\Theta_{0})=1$, $\pi_{B}(\theta\in\Theta_{0})=0$, and $\pi_{A}(Y=y)=\pi_{B}(Y=y)$ for all $y\in\{0,\ldots,n_{1}\}\times\{0,\ldots,n_{0}\}$.
Proof.
Define $\tilde{\pi}_{A}=\tilde{\pi}_{(0.5,0.5,0,0)}$ and $\tilde{\pi}_{B}=\tilde{\pi}_{(0,0,0.5,0.5)}$. The prior $\tilde{\pi}_{A}$ arises from assuming that the units in the finite population were sampled i.i.d. from a superpopulation in which half of individuals are always-takers and the other half are never-takers. The prior $\tilde{\pi}_{B}$ arises analogously if the superpopulation is half compliers and half defiers. As argued in the proof to Proposition˜3.1, the marginal distribution of $Y$ after (i) sampling $\theta\sim\mathrm{Multinomial}(1,p)$ and (ii) generating $Y\sim f_{\theta}$ depends on $p$ only through $p_{at}+p_{c}$ and $p_{at}+p_{d}$. It follows that the two priors imply the same unconditional distribution for $Y$, $\tilde{\pi}_{A}(Y)=\tilde{\pi}_{B}(Y)$ for all $Y$.
Now, consider type counts $\theta_{0}(n_{at})=(n_{at},n-n_{at},0,0)$, which have positive numbers of only always-takers and never-takers. Observe that the support of $Y$ under $\theta_{0}$ is all points of the form $(m,n_{at}-m)$ for $m=\max\{0,n_{at}-n_{0}\},\ldots,\min\{n_{1},n_{at}\}$, where $(m,n_{at}-m)$ is the realization of $Y$ if $m$ of the always-takers are selected for treatment. Hence the support points for $Y$ under $\theta_{0}$ lie on a line with slope of 1-1 and intercept $n_{at}$. Letting $S^{Y}_{\theta_{0}}$ denote the support of $Y$ under $\theta_{0}$, we see that the $S^{Y}_{\theta_{0}}$ are disjoint for all $\theta_{0}$ to which $\tilde{\pi}_{A}$ assigns positive support:
$S_{\theta_{0}(n_{at})}^{Y}\cap S_{\theta_{0}(n_{at}^{\prime})}^{Y}=\varnothing\text{ for $n_{at}\neq n_{at}^{\prime}$}.$
$\tilde{\pi}_{A}(Y=y\mid\theta\in\tilde{\Theta}_{0})=\tilde{\pi}_{A}(Y=y\mid Y\in S^{Y}_{\tilde{\Theta}_{0}}),$
Similarly, let $\theta_{1}(n_{c})=(0,0,n-n_{c},n_{c})$, which has only compliers and defiers. The support points of $Y$ under $\theta_{1}$ take the form $(m,n_{0}-(n_{c}-m))=(m,m+n_{0}-n_{c})$ for $m=\max\{0,n_{c}-n_{0}\},\ldots,\min\{n_{1},n_{c}\}$, corresponding to the case where $m$ compliers are selected for treatment. These points lie on an upward sloping line with slope 1 and intercept $n_{0}-n_{c}$. Letting $S^{Y}_{\theta_{1}}$ denote the support of $Y$ under $\theta_{1}$, we again see that the $S^{Y}_{\theta_{1}}$ are disjoint. Hence, for any set $\tilde{\Theta}_{1}$ of the form $\{\theta_{1}(n_{c,1}),\ldots,\theta_{1}(n_{c,K})\}$, we have that
$\tilde{\pi}_{B}(Y=y\mid\theta\in\tilde{\Theta}_{1})=\tilde{\pi}_{B}(Y=y\mid Y\in S^{Y}_{\tilde{\Theta}_{1}}),$
Now, let
$\tilde{\Theta}_{0}=\begin{cases}\{\theta_{0}(n_{at})\mid n_{at}=0,2,...,n\}&n_{1}\text{ is odd}\\ \{\theta_{0}(n_{at})\mid n_{at}=1,3,...,n-1\}&n_{1}\text{ is even}\end{cases}\subset\Theta_{0}$
and $\tilde{\Theta}_{1}=\{\theta_{1}(n_{c})\mid n_{d}=1,3,\ldots,n-1\}\subset\Theta_{1}$. Lemma˜A.2 shows formally that
$S^{Y}_{\tilde{\Theta}_{1}}=S^{Y}_{\tilde{\Theta}_{0}}.$
An intuitive illustration of the proof is shown in Figure˜2 for the setting where $n=4$, in which case $S_{\tilde{\Theta}_{1}}$ corresponds to the union of the points on the two upward-sloping lines in orange, and $S_{\tilde{\Theta}_{0}}$ corresponds the unions of the points on the two downward-sloping lines in blue.
[…]
Figure 2. Illustration of the sets $S_{\tilde{\Theta}_{0}}$ and $S_{\tilde{\Theta}_{1}}$
$\pi_{A}(Y=y)=\tilde{\pi}_{A}(Y=y\mid Y\in S^{Y}_{\tilde{\Theta}_{0}})=\tilde{\pi}_{B}(Y=y\mid Y\in S^{Y}_{\tilde{\Theta}_{1}})=\pi_{B}(Y=y).$
This completes the proof.
Lemma A.2.
Fix an even $n$ and let $n_{1}=n_{0}=n/2$. Let
$B=\left\{{(m,n_{0}-(n_{c}-m)):m=\max(0,n_{c}-n_{0}),\ldots,\min(n_{1},n_{c}),n_{c}=1,3,5,\ldots,n-1}\right\}$
and
$A=\left\{{(m,n_{a}-m):m=\max(0,n_{a}-n_{0}),\ldots,\min(n_{1},n_{a}),n_{a}=\begin{cases}1,3,\ldots,n-1&n/2\text{ is even}\\ 0,2,\ldots,n&n/2\text{ is odd}\end{cases}}\right\}.$
Then
[…]
Proof.
It is easy to see that $A$ and $B$ are both subsets of $n/2]\times[n/2]$. Observe that $B$ consists of opposite-parity $(a,b)$ if $n_{0}$ is even and consists of same-parity points if $n_{0}$ is odd. Therefore $B\subset C$. Similarly, $A$ consists of opposite-parity points if $n/2$ is even and same-parity points if $n/2$ is odd. Therefore $A\subset C$.
Now, suppose $n/2$ is even and consider $(a,b)\in[n/2]\times[n/2]$ of opposite parity. Let $n_{a}=a+b$ and $m=a$. Then $(m,n_{a}-m)\in A$, since $n_{a}$ is odd and bounded by $n$. Similarly, let $m=a$ and $n_{c}=m+n_{0}-b$. Since $n_{c}$ is odd and $0\leq n_{c}\leq n$, we conclude that $(a,b)\in B$ as well. This proves that $A=B=C$ when $n/2$ is even. The proof for $A=B=C$ when $n/2$ is odd is analogous. ∎
See 4.2
Proof.
By Lemma˜A.1, there exists priors $\pi_{A}$ and $\pi_{B}$ such that $\pi_{A}(\theta\in\Theta_{0})=1$, $\pi_{B}(\theta\in\Theta_{1})$, and $\pi_{A}(Y=y)=\pi_{B}(Y=y)$ for all $y$. Observe that
and hence
It follows that there exists some $\theta_{1}\in\operatorname{supp}(\pi_{B})\subseteq\Theta_{1}$ such that $E_{\theta}[\delta(Y)]\leq\alpha$, which gives the desired result. ∎
@@
Suppose, without loss of generality, that $n_{1}\leq n_{0}$ (if not, we can adopt the same argument reversing the roles of $n_{d}$ and $n_{c}$). Let $\theta_{1}$ be the type with $n_{d}=n_{1}-1$, $n_{c}=n-n_{d}$, and $n_{at}=n_{nt}=0$. Note that $n_{c}=n-n_{1}+1\geq 2$. Observe that under $\theta=\theta_{1}$, the support of $Y$ is
@@
Now, for any $\theta_{0}\in\Theta_{0}$, let $S^{Y}_{\theta_{0}}$ denote the support of $Y$ under $\theta=\theta_{0}$. We claim that $S^{Y}_{\theta_{0}}\not\subseteq S^{Y}_{\theta_{1}}$ for all $\theta_{0}\in\Theta_{0}$.
@@ 1. Case 1
    Suppose that $\theta_{0}$ corresponds to there only being one type in the population. Then $S_{\theta_{0}}^{Y}$ is a singleton set with element either $(0,0)$, $(n_{1},0)$, $(n_{1},n_{0})$ depending on whether the lone type is NT, C, or AT. It is clear that $(0,0)\not\in S_{\theta_{1}}^{Y}$. Next, note that the only possible element of $S_{\theta_{1}}^{Y}$ with first element equal to $n_{1}$ is $(n_{1},n_{d})$, corresponding to the case where $m=n_{1}=\min\{n_{1},n_{c}\}$. However, by construction we have that $0<n_{d}=n_{1}-1<n_{0}$, and thus $(n_{1},n_{d})\neq(n_{1},0)$ and $(n_{1},n_{d})\neq(n_{1},n_{0})$.
@@ 2. Case 2
    Since the treatment effects $(y_{k}(1)-y_{k}(0))$ and $(y_{j}(1)-y_{j}(0))$ are each in $\{-1,0,1\}$, their difference can equal an even non-zero integer only if one is equal to 1 and the other is equal to 1-1. However, this contradicts $\theta_{0}$ satisfying monotonicity. It follows that $S^{Y}_{\theta_{1}}\not\subseteq S^{Y}_{\theta_{0}}$, as we wished to show.

### v2

@@ 5. Bayesian updating
    Some Bayesians update: there exists a prior $\pi$ with $\pi(\theta\in\Theta_{0})\in(0,1)$ such that $\pi(\theta\in\Theta_{0}\mid Y)\neq\pi(\theta\in\Theta_{0})$ with positive $\pi$-probability.
@@
    There exist (nontrivial) Bayesian priors over $\theta$ that never update about the probability that monotonicity holds: if $n$ is even, then for any $c\in(0,1)$, there exists a prior distribution $\pi$ over $\theta$ such that $\pi(\theta\in\Theta_{0}\mid Y)=\pi(\theta\in\Theta_{0})=c$, $\pi$-almost surely.
@@
The existence of a prior that does not update also implies that any binary classifier of whether $\theta\in\Theta_{0}$ does no better than random guessing for some parameter values $\theta$. Specifically, given $a\in\{0,1\}$, let $\ell(a,\theta)=1\{a\neq 1\{\theta\in\Theta_{0}\}\}$ be an indicator for whether $a$ mis-classifies whether $\theta\in\Theta_{0}$ (i.e. zero-one classification loss). Let $c:\mathcal{Y}\to[0,1]$ be a possibly-randomized classifier that sets $a=1$ with probability $c(Y)$, and let $R(c,\theta)=E_{\theta}[c(Y)\ell(1,\theta)+(1-c(Y)\ell(0,\theta)]$ be its average classification error at parameter $\theta$ (risk under 0-1 loss).
Corollary 5.1.
Suppose $n$ is even. Then $\inf_{c(\cdot)}\sup_{\theta}R(c,\theta)=0.5$.
Note that a trivial classifier that guesses randomly ($c(y)=0.5,$ $\forall y$) achieves a 0.5 mis-classification rate. Corollary˜5.1 implies that every classifier does no better than this at some parameter value $\theta$ (when $n$ is even). This follows from Part 2 of Proposition˜5.1: there exists a prior that never updates about whether $\theta\in\Theta_{0}$, and hence Bayes risk under this prior must be trivial. christy_counting_2025 consider the maximum likelihood estimator $\hat{\theta}$ and say that it “provides evidence in favor of monotonicity” if $\max\{\hat{\theta}_{c},\hat{\theta}_{d}\}=0$. Corollary˜5.1 implies that this classification of whether the data support monotonicity does no better than random guessing at some parameter values. [^11] In fact, in some cases, this classification rule does worse than random guessing. Consider Example 1 above with $n=2$. Note that with 1 complier and 1 defier, the support of $Y$ is $\{(0,0),(1,1)\}$, with each support point occurring with probability $1/2$. However, $Y=(1,1)$ with probability 1 if there are 2 always-takers, and $Y=(0,0)$ with probability 1 if there are two never-takers. Thus, the MLE never corresponds to having 1 complier and 1 defier, so worst-case misclassification error is 1. Moreover, we cannot “fix” this undesirable property by choosing a different classification rule.
Part 2 of Proposition˜5.1 and Corollary˜5.1 focused on the case where $n$ is even. In the appendix, we show that when $n$ is odd, there exist non-degenerate priors that minimally update in the sense that the expected change between the prior and posterior is small: Proposition˜A.1 shows that there exists a prior such that
$E_{Y\sim\pi_{Y}}[|\pi(\theta\in\Theta_{0}\mid Y)-\pi(\theta\in\Theta_{0})|]=O(2^{-n}),$
where $\pi_{Y}$ is the distribution of $Y$ induced by prior $\pi$. This implies Corollary˜A.1, which states that when $n$ is odd, a classifier of whether $\theta\in\Theta_{0}$ has worst-case mis-classification error at least $0.5-O(2^{-n})$.
@@ 6. Conclusion
We study what a completely randomized experiment can reveal about finite-population monotonicity. We show that from the design-based perspective, the type counts $\theta$ in the finite population are in fact identified. However, the extent to which we can feasibly learn about violations of monotonicity is severely limited: frequentist tests generically have poor power, and some Bayesians never update about whether the null is true. Thus, formal identification translates to little, if any, practical learning about monotonicity. These results highlight that conclusions about identification may differ depending on whether one adopts a sampling-based versus design-based perspective, and that studying the properties of frequentist tests and Bayesian updating may provide a more realistic assessment of the extent to which learning is possible in design-based settings. An interesting avenue for future research is to explore whether similar issues arise in other design-based causal inference problems.
@@ Appendix A Proofs
For the first claim, it suffices to construct a mapping from $f_{\theta}$ to $\theta$. Fix $q=n_{1}/n$ as the treatment probability. Observe that, by linearity of expectation,
@@
Second, label underlying units $i=1,\ldots,n$ and collect always takers in $N_{at}\subset[n]$, compliers in $N_{c}\subset[n]$, and defiers in $N_{d}\subset[n]$. Let $N_{1}\subset[n]$ collect the (random) set of treated individuals. Observe that
$E_{\theta}[Y_{T}Y_{U}]$ $=\sum_{i\neq j,i\in N_{at}\cup N_{c},j\in N_{at}\cup N_{d}}P(\text{$i\in N_{1}$, $j\not\in N_{1}$})$ (Linearity of expectations)
$=|\{(i,j)\,:\,i\neq j,i\in N_{at}\cup N_{c},j\in N_{at}\cup N_{d}\}|\cdot P(\text{$1\in N_{1}$, $2\not\in N_{1}$})$ (Symmetry)
$=\left[{(n_{at}+n_{c})(n_{at}+n_{d})-n_{at}}\right]P(\text{$1\in N_{1}$, $2\not\in N_{1}$}).$
Thus, for $r=P(\text{$1\in N_{1}$, $2\not\in N_{1}$})=\frac{n_{1}n_{0}}{n(n-1)}$,
$(n_{at}+n_{c})(n_{at}+n_{d})-n_{at}=\frac{E_{\theta}[Y_{T}Y_{U}]}{r)}$
is a functional of $f_{\theta}$. Letting $(\mu_{T},\mu_{U},\mu_{TU}):=(E_{\theta}[Y_{T}],E_{\theta}[Y_{U}],E_{\theta}[Y_{T}Y_{U}])$, we see that
$\frac{\mu_{T}}{q}=n_{at}+n_{c},\hskip 28.45274pt\frac{\mu_{U}}{1-q}=n_{at}+n_{d},\hskip 28.45274pt\frac{\mu_{TU}}{r}=\frac{\mu_{T}}{q}\,\frac{\mu_{U}}{1-q}-n_{at}.$
@@
$n_{at}=\frac{\mu_{T}}{q}\,\frac{\mu_{U}}{1-q}-\frac{\mu_{TU}}{r},\hskip 28.45274ptn_{c}=\frac{\mu_{T}}{q}-n_{at},\hskip 28.45274ptn_{d}=\frac{\mu_{U}}{1-q}-n_{at}.$
Since $(\mu_{T},\mu_{U},\mu_{TU})$ is a function of $f_{\theta}$, we have therefore shown that there is an inverse mapping from $f_{\theta}$ to $\theta$, and hence the mapping $\theta\mapsto f_{\theta}$ must be one-to-one, as needed.
@@
Suppose, without loss of generality, that $n_{1}\leq n_{0}$ (if not, we can adopt the same argument reversing the roles of $n_{d}$ and $n_{c}$). Let $\theta_{1}$ be the type with $n_{d}=n_{1}-1$, $n_{c}=n-n_{d}$, and $n_{at}=n_{nt}=0$. Note that $n_{c}=n-n_{1}+1\geq 2$. Observe that the support of $Y$ under $\theta_{1}$ is
@@
Now, for any $\theta_{0}\in\Theta_{0}$, let $S^{Y}_{\theta_{0}}$ denote the support of $Y$ under $\theta_{0}$. We claim that $S^{Y}_{\theta_{0}}\not\subseteq S^{Y}_{\theta_{1}}$ for all $\theta_{0}\in\Theta_{0}$.
@@ 1. Case 1
    Suppose that $\theta_{0}$ corresponds to there only being one type in the population. Then $S_{\theta_{0}}^{Y}$ is a singleton set with element either $(0,0)$, $(n_{1},0)$, $(0,n_{0})$ $(n_{1},n_{0})$ depending on whether the lone type is $nt$, $c$, $d$ or $at$. It is clear that $(0,0)\not\in S_{\theta_{1}}^{Y}$ and $(0,n_{0})\not\in S_{\theta_{1}}^{Y}$ since the first coordinate of points in $S_{\theta_{1}}^{Y}$ is strictly positive by construction. Next, note that the only possible element of $S_{\theta_{1}}^{Y}$ with first element equal to $n_{1}$ is $(n_{1},n_{d})$, corresponding to the case where $m=n_{1}=\min\{n_{1},n_{c}\}$. However, by construction we have that $0<n_{d}=n_{1}-1<n_{0}$, and thus $(n_{1},n_{d})\neq(n_{1},0)$ and $(n_{1},n_{d})\neq(n_{1},n_{0})$.
@@ 2. Case 2
    Since the treatment effects $(y_{k}(1)-y_{k}(0))$ and $(y_{j}(1)-y_{j}(0))$ are each in $\{-1,0,1\}$, their difference can equal an even non-zero integer only if one is equal to 1 and the other is equal to 1-1. However, this contradicts $\theta_{0}$ satisfying monotonicity. It follows that $S^{Y}_{\theta_{0}}\not\subseteq S^{Y}_{\theta_{1}}$, as we wished to show.